Information:

Using the money from their recent bond issue, Terry’s management has decided to declare an additional $562,500 dividend. The date of declaration is December 30, Year 3. The date of record will be January 15, Year 4, and the date of payment will be January 30, Year 4.                                             As an additional signal to the market, Terry’s management repurchased 205,000 shares of Terry’s common stock on December 15, Year 3 for $8.00 a share. Tax rate is 25%

Terry’s management would like to know the effect of the sale on the following ratios:

-Current Ratio                                                                           -ROA

Assignment:

1.Calculate each of the two (2) ratios Before you make any adjustments.

2.Make the appropriate journal entries, if any, to account for Terry’s extra dividend and stock repurchase (including any necessary changes to income tax expense).

3.Make any necessary changes to the financial statements.

4. At the beginning of Year 2, Terry’s Board of Directors authorized stock options for the executive team. The team could buy shares from the company at a set price for $10.00/share, then sell them at the current market price to make a bonus. On December 10th, Terry’s stock price was only $8.40/share, making these stock options worthless. Based on this information, do you think that the management teams’ decision to authorize an additional dividend and repurchase shares on December 15th was ethical? Why or why not?

Terry Co. Multi-Step Income Statement For the Year Ended December 31, Year3 566,600,000 $5,827,500 $6,493,500 560,106,500 S23,572,487 Miscellaneous Selling Expenses Expense S6,930,563 Depreciation & Amortization Expense $1,998,000 Miscellaneous Admin. Expenses $499,500 Total Administrative Expenses S6,942,634 S13,873,197 $9,699,290 S208,125 57 Income from Continuing Operations before Taxes 9,450,361 362 57.087.771 $2.66

Terry Co. Balance Sheet As of 12/31/Year 3 Year 2 Current Assets Cash A/R Allowance for Bad Debts Inventory $5,554,969 $3,330,000 $5,994,000 $5,661,000 (S333,000) (1,665,000) $7,992,000 $9,324,000 S499,500 $999,000 $832.500$666000 $20,539.969 $18,315,000 Prepaid Utilities Total Current Assets Long-term Investments Loans to other businesses Expansion Fund $2,664,000 $2,664,000 $2,843,560 $2,843,560 $5,507,560 $5,507,560 Total Long-term Investments PPE Land Building $7,326,000 $4,662,000 $5,328,000 $5,328,000 $18,648,000 $8,658,000 Accumulated Depreciation Total PPE $22,644,000 $11,988,000 Intangible Assets Patents, net S999.000 $49,690,529 $36,809,560 Total Assets Liabilities and Stockholders Equity Current Liabilities Accounts Payable Income Tax Payable Interest Payable Unearned Revenue Wages Payable Current Portion of Loan Payable $3,044,570 $3,996,000 $2,322,880 $666,000 S0 1,198,800 $999,000 $666,000 $832.500 $333,000 S333.000 $7,595,250 $6,826,500 S30,000 Total Current Liabilities Long-term Debt Loan Payable Bonds Payable, net Notes Payable $3,663,000 $3,996,000 SO $9,324.000 $5,328,000 $14,548,447 $9,324,000 $22,143,698 $16,150,500 $1,561,447 Total Long-term Debt Total Liabilities Stockholders Equity Common stock 2,660,000 2,660,000 ($1 par, 4,655,000 authorized, 2,660,000 outstanding) Additional Paid-In capital Retained Eamings Accumulated OCI $1,998,000 $1,998,000 $23,633,803 $16,746,032 Total Stockholders Equity Total Liabilities and Stockholders Equity $27,546,831 $20,659,060 $49,690,529 $36,809,560

Terry Co. Statement of Cash Flows For Year Ended 12/31/Year 3 Cash Flow from Operations Net Income $7.087.771 Change in A/R Change in Inventory Change in Prepaid Insurance Change in Prepaid Utilities ($1,665,000) S1,332,000 $499,500 ($166,500) $1,998,000 $2,479 ($951,430) 1,656,880 $30,000 $199,800 Amortization of Bond Discount Change in A/P Change in Income Tax Payable Change in Interest Payable Change in Uneamed Revenue Change in Wages Payable Net Cash Flow from Operations S9,857,000 Cash Flow from Investments Purchase of Land Purchase of Equipment ($2,664,000) Net Cash Flow from Investments ($12,654,000) Cash Flow from Financing Repayment of Loans Issuance of Bonds Payable Issuance of Notes Payable Payments of Dividends ($333,000) $1,558,969 S3,996,000 Net Cash Flow from Financing 1,969 Net Increase (Decrease) in Cash Cash, January 1, Year 2 Cash, December 31, Year 2 $2,224,969 S3,330.000 $5,554,969
Answer #1
2..Year 3
15-Dec Treasury stock 1640000
Cash 1640000
(205000*8)
30-Dec Retained earnings 562500
Dividends payable 562500
Dividends are paid out of after-tax income & hence, no change is necessary to the income-tax expense a/c.
1.Before adjustment 3.After adjustment
Current Ratio=Current Assets/Current Liabilities
20539969/7595250= (20539969-1640000)/(7595250+562500)=
2.70 2.32
Return on Assets(ROA)=Net Income/Total assets
7087771/49690529= 7087771/(49690529-1640000)=
14.26% 14.75%
Balance sheet (after adjustment)
Items(Numbers) that will change
All others remain same
Assets
Cash 5554969-1640000= 3914969
Total Current assets 20539969-1640000= 18899969
Total assets 49690529-1640000= 48050529
Liabilities & SH Equity
Current liabilities
Dividend payable 562500 562500
Total current Liabilities 7595250+562500= 8157750
Stockholders' equity
Retained earnings 23633803-562500= 23071303
Treasury stock -1640000 -1640000
Total Stockholders' equity 27546831-562500-1640000= 25344331
Total Liabilities & SH Equity 49690529-1640000= 48050529


4..There is nothing unethical about the management teams’ decision to authorize an additional dividend and repurchase shares ,when the executive stock options are pending to be exercised for the following reasons:
Normally when stock options are exercised,it leads to more number of shares outstanding & hence the earning per share(EPS) goes down and also has its effect on market price & the P/E ratio--which in turn affects the stock's marketability.
So, to obviate all such consequences, companies buy back some of its shares ,to maintain the market price of its stock, by not allowing the EPS to be diluted.
Secondly, the share price in the market has come down to $ 8.40 /share--so executives can buy the share from the market itself at $ 8.40 instead of exercising their option at $ 10 /share with the company.
The options,in any case , have become unexercisable , due to market forces.